c) Determine the location and values of the absolute maximum and absolute minimum for the given function: 𝑓(π‘₯) = (βˆ’π‘₯ + 2) ΰ¬Έ , π‘€β„Žπ‘’π‘Ÿπ‘’ 0 ≀ π‘₯ ≀ 3


1
Expert's answer
2021-06-20T18:31:17-0400
"F(x)=(-x+2)^3"

"0\leq x\leq 3"

Find the first derivative



"F'(x)=((-x+2)^3)'=-3(-x+2)^2"

Find the critical number(s)



"F'(x)=0=>-3(-x+2)^2=0""x=2"



Differentiate f'(x) w.r.t. x-

F''(x)=6(-x+2)

F''(2)=6(-2+2)=0


So F(x) has neither maximum nor minimum at x=2.


"F(0)=(-0+2)^3=8\\

F(3)=(-3+2)^3=-1\\

F(2)=(-2+2)^3=0"


So F(x) has absolute maximum at x=0 and The value is 8.


F(x) has absolute minimum at x=3 and The value is -1.